How many moles of O2 can be produced by letting 6.0 moles of KClO3 react based on the below equation?

2KClO sub 3, reaction arrow, 2KCl + 3O sub 2.
1
Expert's answer
2015-05-25T11:01:28-0400
The equation is:
2KCIO3 = 3KCI + 3O2
The molar ratio of KClO3 to O2 in this reaction is 2/3
Using this ratio we can find the number of moles of O2: 2/3=6/x; x=(3 x 6)/2 = 9 moles